OpenStudy (anonymous):

1. m + 5 = –7.3 2. m + 5 = 7.3 3. m + 7.3 = –5 4. m + 7.3 = 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5

OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 1. Which related equation can be used to solve the equation? Column A Column B 1. m + 5 = –7.3 2. m + 5 = 7.3 3. m + 7.3 = –5 4. m + 7.3 = 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5